# GarmentCodeData: A Dataset of 3D Made-to-Measure Garments With Sewing Patterns

### v1

@@ Abstract
The dataset will be released with or before the official publication of this work. The code will be shared publicly, subject to licenses of third-party components.
@@ 3.1 Overview
The data generation process consists of three steps: body shape sampling, sewing pattern sampling, and draping. We sample 5,000 body shapes from a custom statistical body model,
and then automatically evaluate the body measurements for each sample with a collection of algorithms designed for different types of body measurements (Sec. 3.3). Apart from body shape variations, we include small variations in the arm angles to increase the robustness of the models trained on our dataset to slight variations in A-poses in the inputs. Next, sewing pattern designs are sampled (Sec. 3.2). We carefully adjust the design space to ensure a diverse yet reasonable design distribution in the final dataset. Every design sample is fitted to the canonical neutral body model (statistical average across genders) and to one of the randomly sampled body shapes based on the evaluated measurements, which provides a ground truth design reference in canonical shape for every sample. The generated sewing patterns are then draped on top of their corresponding body models using our custom draping pipeline, implemented using Warp [22], a GPU-based XPBD simulator.
@@ 3.2 Design space preparation and sampling
Secondly, we notice that usually, most of the design elements in daily wear follow a common style convention, with only some elements varying in style to add uniqueness to a particular piece. We implement this idea by introducing sampling probabilities to the default value of each design parameter . We flip a coin with a probability to draw the default value of . If the default value is not selected, we draw a non-default value from the corresponding parameter range uniformly. The default values and their probabilities are defined manually to ensure reasonable and interesting designs and more common features to appear more often in the data samples. By adjusting the default value probabilities for the top-level garment structure, we set up roughly a third of the samples to be upper-body garments, another third – lower-body garments, and the rest to be full-body garments (dresses and jumpsuits). We define a strong preference for symmetric garments by setting for the parameter that controls left-right asymmetry in necklines and sleeves.
@@ 4.1 Body measurements quality
The measurements are on average overestimated by for the bust circumference, for the hip circumference and for the leg circumference. High errors arise due to possible tangential drift (bust measurement) and when the scanned individuals are wearing wide, non-skin-tight pants (leg circumference). Overestimations in our measurements may also arise due to the missing compression of soft tissue when measuring a static mesh. The shoulder width is underestimated on average by our method, as we measure on top of the shoulder ball joint, where the garment shoulder line is typically located. The ground truth measurement is taken on the left and right sides of the deltoid muscles. The overestimation of the circumference measurements does not have a negative impact on the fitting of the garments in our experiments.
@@ 5 Discussion and limitations
We presented a generation pipeline for sampling 3D made-to-measure garments with ground-truth sewing patterns, as well as GarmentCodeData – a dataset of 115,000 3D garments of different designs fitted to a variety of body shapes. Although we aimed to create a large-scale dataset suitable for a sewing pattern reconstruction task, we envision other potential applications of the pipeline. The proposed probability-based design space control allows for its easy adjustment and creation of specialized sets of designs, for example, specific garment types (workwear, uniforms, etc.). The availability of various designs fits to a neutral body shape in the dataset opens ups applications for design retargeting. Using the pipeline for obtaining references for neural cloth simulation research could also be an interesting direction to explore.
@@
Our work improves on some of the limitations of existing works, but there is still ample room for creating truly comprehensive datasets. First, we inherit some limitations of the GarmentCode [14] patterns, addressing which would require additional effort put into the draping pipeline: lack of fabric layers, sharp pleats, hardwear (e.g., zippers, buttons or hooks), the ability to represent full outfits. Our take on fabric material distribution is rather minimalistic, and further effort is needed to model materials according to real-world fabric properties. As discussed in Sec. 4.2, some of the simulation errors we encounter are due to issues with initial body-pattern alignment and methods for body-cloth collision resolution, which require further research. As for body measurements, our approach may result in inaccurate measurements due to errors in surface fit, especially on extreme body shapes. Such errors may lead to the topology-based initial guess being far off, and the correct horizontal plane may not be found within the given optimization bounds.
@@
Our work required a sizable amount of computational power to generate large enough amounts of data and run experiments. However, the server infrastructure we use is carbon neutral, and is running fully on renewable energy [reference omitted for anonymity], which minimizes the climate impact of this work.
@@
This work was partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (ERC Consolidator Grant, agreement No. 101003104, MYCLOTH). Stephan Wenninger has been funded by “Stiftung Innovation in der Hochschullehre” through the project “Hybrid Learning Center” (FBM2020-EA-690-01130). We thank Ami Beuret for his help in refactoring GarmentCode codebase. We are grateful to Jana Schuricht for her professional consultations on patternmaking and to members of IGL and GGG for their continuous support and cheer throughout this project.

### v3

@@ Abstract
The links to the dataset and available code are shared on the project page.111https://igl.ethz.ch/projects/GarmentCodeData/
@@ 3.1 Overview
The data generation process consists of three steps: body shape sampling, sewing pattern sampling, and draping. We sample 5,000 body shapes from a custom statistical body model, and then automatically evaluate the body measurements for each sample with a collection of algorithms designed for different types of body measurements (Sec. 3.3). Apart from body shape variations, we include small variations in the arm angles to increase the robustness of the models trained on our dataset to slight variations in A-poses in the inputs. Next, sewing pattern designs are sampled (Sec. 3.2). We carefully adjust the design space to ensure a diverse yet reasonable design distribution in the final dataset. Every design sample is fitted to the canonical neutral body model (statistical average across genders) and to one of the randomly sampled body shapes based on the evaluated measurements, which provides a ground truth design reference in canonical shape for every sample. The generated sewing patterns are then draped on top of their corresponding body models using our custom draping pipeline, implemented using Warp [22], a GPU-based XPBD simulator.
@@ 3.2 Design space preparation and sampling
Secondly, we notice that usually, most of the design elements in daily wear follow a common style convention, with only some elements varying in style to add uniqueness to a particular piece. We implement this idea by introducing sampling probabilities to the default value of each design parameter . We flip a coin with a probability to draw the default value of . If the default value is not selected, we draw a non-default value from the corresponding parameter range uniformly.
The default values and their probabilities are defined manually to ensure reasonable and interesting designs and more common features to appear more often in the data samples. By adjusting the default value probabilities for the top-level garment structure, we set up roughly a third of the samples to be upper-body garments, another third – lower-body garments, and the rest to be full-body garments (dresses and jumpsuits). We define a strong preference for symmetric garments by setting for the parameter that controls left-right asymmetry in necklines and sleeves.
@@ 4.1 Body measurements quality
The measurements are on average overestimated by for the bust circumference, for the hip circumference and for the leg circumference.
High errors arise due to possible tangential drift (bust measurement) and when the scanned individuals are wearing wide, non-skin-tight pants (leg circumference).
Overestimations in our measurements may also arise due to the missing compression of soft tissue when measuring a static mesh. The shoulder width is underestimated on average by our method, as we measure on top of the shoulder ball joint, where the garment shoulder line is typically located. The ground truth measurement is taken on the left and right sides of the deltoid muscles.
The overestimation of the circumference measurements does not have a negative impact on the fitting of the garments in our experiments.
@@ 5 Discussion and limitations
We presented a generation pipeline for sampling 3D made-to-measure garments with ground-truth sewing patterns, as well as GarmentCodeData – a dataset of 115,000 3D garments of different designs fitted to a variety of body shapes. Although we aimed to create a large-scale dataset suitable for a sewing pattern reconstruction task, we envision other potential applications. The proposed probability-based design space control allows for its easy adjustment and creation of specialized design sets, for example, specific garment types (workwear, uniforms, etc.). The availability of various designs fits to a neutral body shape in the dataset opens up applications for design retargeting. Using the pipeline for neural cloth simulation research could also be an interesting direction to explore.
@@
Our work improves on some of the limitations of existing works, but there is still ample room for creating truly comprehensive datasets. First, we inherit some limitations of the GarmentCode [14] patterns, addressing which would require additional effort put into the draping pipeline: lack of fabric layers, sharp pleats, hardwear (e.g., zippers, buttons or hooks), the ability to represent full outfits. Our take on fabric material distribution is rather minimalistic, and further effort is needed to model materials according to real-world fabric properties. As discussed in Sec. 4.2, some of the simulation errors we encounter are due to issues with initial body-pattern alignment and methods for body-cloth collision resolution, which require further research.
As for body measurements, our approach may result in inaccurate measurements due to errors in surface fit, especially on extreme body shapes. Such errors may lead to the topology-based initial guess being far off, and the correct horizontal plane may not be found within the given optimization bounds.
@@
Our work required a sizable amount of computational power to generate the data. However, the server infrastructure we use is carbon neutral, and is running fully on renewable energy [32], which minimizes the climate impact of this work.
@@
This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (ERC Consolidator Grant, agreement No. 101003104, MYCLOTH). Stephan Wenninger has been funded by “Stiftung Innovation in der Hochschullehre” through the project “Hybrid Learning Center” (FBM2020-EA-690-01130). We thank Ami Beuret for his help in refactoring GarmentCode codebase. We are grateful to Jana Schuricht for her professional consultations on patternmaking and to members of IGL and GGG for their continuous support and cheer throughout this project.